# The Semantic Hub Hypothesis: Language Models Share Semantic Representations Across Languages and Modalities

### v1

@@ 1 Introduction
To what extent is this idealized strategy adopted by actual models? Wendler et al. (2024) find that on simple synthetic tasks, Llama-2 (Touvron et al., 2023) maps various input languages into a shared “English space” before projecting back out into another language, hinting that it leverages this shared representation scheme to an extent, at least for different languages. We show that this is a much more general phenomenon: when a model processes inputs from multiple data types, there is a shared representation space, and this space is scaffolded by the LM’s inherently dominant data type (usually English). By scaffolded, we mean that the shared space can be interpreted to an extent in the dominant data type via the logit lens (nostalgebraist, 2020). Following the cognitive science nomenclature, we call this shared representation space the LM’s “semantic hub.”
@@
Our work is complementary and distinct from prior work which finds structural similarities between the representation spaces of models trained (usually independently) on different data types, such as those showing that text representations from text-only LMs can be aligned to vision/audio representations of modality-specific models (Ilharco et al., 2021; Merullo et al., 2022; Li et al., 2023; Ngo & Kim, 2024; Huh et al., 2024; i.a.), and the literature on cross-lingual word embedding alignment (Mikolov et al., 2013; Artetxe et al., 2017; Conneau et al., 2018; Schuster et al., 2019; i.a.). We instead show that an LM trained on multiple data types represents and processes them in a shared space without requiring explicit alignment. We hope our findings shed light on ways to more easily interpret the mechanisms of current models and motivate future work aimed at better controlling models using these insights. [^2] We release our code at https://github.com/ZhaofengWu/semantic-hub.
@@ 2 The Semantic Hub Hypothesis
An LM parameterizes a similar process: it uses $M_{\text{LM}}$ to map various input data types into a representation space $\mathcal{S}_{\text{LM}}\subseteq\mathbb{R}^{d}$ (early layers), performs computations in the space (middle layers), and verbalizes the output via $V_{\text{LM}}$ (end layers and the LM head). However, it is unknown as to how different data types are structured in the representation space. For example, one possibility is that the LM partitions $\mathbb{R}^{d}$ into disjoint subspaces for each data type and processes them separately. We instead hypothesize that LMs, through training, learn to represent and process different data types in a shared representation space that functions as a modality-agnostic “semantic hub.” That is, semantically similar inputs $w_{1:t}^{z_{1}}$ and $w_{1:t^{\prime}}^{z_{2}}$ from distinct data types—for example texts in different languages that are mutual translations—are similarly mapped in $\mathcal{S}_{\text{LM}}$; informally, $M_{\text{LM}}(w_{1:t}^{z_{1}})\approx M_{\text{LM}}(w_{1:t^{\prime}}^{z_{2}})$. However, absolute similarity measures (i.e., $\operatorname{sim}(M_{\text{LM}}(w_{1:t}^{z_{1}}),M_{\text{LM}}(w_{1:t^{\prime}}^{z_{2}}))$) are generally difficult and unintuitive to interpret in high dimensional spaces (see for example Beyer et al. (1999)). We thus focus on relative similarity measures, taking a semantically unrelated sequence $u_{1:t^{\prime}}^{z_{2}}$, and evaluating whether $w_{1:t}^{z_{1}}$ is closer to $w_{1:t^{\prime}}^{z_{2}}$ than $u_{1:t^{\prime\prime}}^{z_{2}}$, as illustrated on the left of Figure 2:
@@ 3.1 Multilingual
Translation datasets enable a direct test for Eq. 1, with semantically equivalent cross-lingual sentences as $w_{1:t}^{z_{1}}$ and $w_{1:t^{\prime}}^{z_{2}}$ and a non-matching sentence as $u_{1:t^{\prime\prime}}^{z_{2}}$. We use the professionally-translated English-Chinese parallel sentences from Chen et al. (2016) ($N=5260$). For each sentence pair, we use a template to transform each sentence and compute the representation cosine similarity for each layer, using the last token position as the sentence representation, which has been shown to preserve information (Morris et al., 2023). We consider two English-dominant LMs, Llama-2 and Llama-3 (Llama-3-Team, 2024), one Chinese-dominant LM, Baichuan-2 (Yang et al., 2023), and one multilingual LM, BLOOM (BigScience, 2023), specifically the 7B/8B variants. See §A.1 for more details.
@@
Figure 4 shows the probability of English and Chinese (and of the top language when it is neither) for each layer on 10,000 English/Chinese Wikipedia sentences. When English-dominant models process Chinese text, Wendler et al.’s finding generalizes, where English dominates in the intermediate layers and Chinese only dominates in the final layers. On the Chinese-dominant LM, this trend flips: when processing English text, its intermediate layers are closer to Chinese space and the final layers are closer to English space. For BLOOM, a multilingual model with a relatively balanced training language mixture, we do not see a clear dominating language in the intermediate layers; when we manually inspect the closest token, in most cases we observe symbols with no clear semantics (though this does not mean it does not have a unified representation space: see Exp. 1). 70B model trends are also highly similar to the 7B/8B ones.
@@ 3.3 Code
Many recent LMs are trained on code corpora (Llama-3-Team, 2024; Gemini-Team, 2024; i.a.). We find that they similarly process code input by projecting them into a unified representation space shared with regular language tokens. Figure 7 shows some examples, where the LM in the intermediate layers tends to verbalize the future in free-form English, unconstrained by program syntax. E.g., in the first program, given the Python prefix “... for idx, e in enumerate(numbers): for idx2, e2 in enumerate(numbers”, instead of the ground-truth continuation in Python “): if idx != idx2:”, the most salient intermediate token is “except”, likely attempting to predict in English “(for each element in numbers) except if it is equal to idx”. Similarly, in a list literal expression “[1.0, 2.0,” instead of continuing in Python “ 3.0”, it predicts “and”, which is a natural way to continue in English. In these cases, it is difficult to obtain semantically equivalent English-Python prefix pairs like in §3.1 for testing Eq. 1444We may argue that functions and their specifications constitute such pairs, but we found that their correspondence is often too abstract and non-exact to manifest as similar representations., so we only test Eq. 3 across a few targeted examples in Python.
@@
Function arguments have names in the definition, such as “range(start, end, step)”; but when invoked, they are filled with actual context-appropriate expressions. We call the argument names “semantic roles”, and the context-specific expressions the “surface forms”, inspired by thematic relations in linguistics (Fillmore, 1968; Jackendoff, 1974; i.a.). With an example in Figure 1, we show that LMs predict the arguments by first “thinking” about their semantic role ($\tau^{z^{\star}}$) and then instantiating with surface-constrained expressions ($\tau^{z^{\circ}}$). We extract all function calls and arguments from MBPP with simple filtering, resulting in 540 arguments (see §A.2 for details). For each argument, we use the logit lens to inspect the hidden states $h_{t}^{\ell}$ at the preceding token (“(” or “,”). For each argument, Figure 8(b) visualizes if the semantic role or the surface form is closer to each layer’s hidden state. The semantic role ($\tau^{z^{\star}}$) dominates for the early to middle layers, and only in the final layers do the representations converge towards the surface form argument ($\tau^{z^{\circ}}$).
@@ 3.5 Visual Input
Though not constituting exact semantic equivalence, an image paired with its caption provides one possible test for Eq. 1. We take 1000 images and corresponding captions in the MSCOCO dataset (Lin et al., 2014) and measure their hidden states cosine similarity in LlaVa-7B (Liu et al., 2023) and Chameleon-7B (Chameleon-Team, 2024). Again, we subtract the average similarity between non-matching image-caption pairs as a baseline, separately for each layer. Figure 11(a) shows that semantically matching inputs across modalities are closer to one another than would be expected from chance, as in the translation experiments.
@@ 4 Intervening in the Semantic Hub
Past work has shown that (monolingual) interventions in the middle layers can steer the output of LMs in predictable ways (Subramani et al., 2022; Turner et al., 2024; Rimsky et al., 2024; i.a.). If the English-dominant LMs have a shared representation space, we should be able to intervene on this space in English even when processing other languages. We use a popular hidden space intervention technique, Activation Addition (ActAdd; Turner et al., 2024), which works by: (1) taking a pair of contrasting steering words that semantically represent the steering effect (e.g., “Good” and “Bad” for sentiment steering), (2) taking their hidden state difference at an intermediate layer, and (3) scaling and adding the steering vector to the hidden states for the original forward pass of the regular generation process, at the same layer, at the beginning of the sequence. We generalize their sentiment-steering experiment cross-lingually. See Turner et al. (2024) for details.
@@
Based on our “semantic role” observation in §3.3, we intervene using the “range” function. We focus on two overloaded versions of “range”: “range(start, end)” and “range(end)”. If the semantic hub causally affects model output, then we can intervene in it to select the function version to use after “range(” in some context, using the English words “start” and “end”. Specifically, we take all single-argument “range(end)” calls in the MBPP dataset ($N=159$) and attempt to expand it into “range(0, end)”. We use a similar intervention method, except simply using the unembeddings of trigger tokens instead of intermediate LM hidden states. We use as trigger tokens (“start” – “end”) and add the intervention vector to the hidden states corresponding to the open parenthesis “(” at an intermediate layer (layer 17). For all these “range” calls in the dataset, we let Llama-2777We do not consider Llama-3 in this case because its default behavior usually generates “range(0, end)” in the first place, and it is unclear how to steer from “range(0, end)” to “range(end)”. generate without and with intervention. Figure 12(b) shows that, with increasing intervention strength, more instances are successfully steered to “range(0, end)”, up to 67%.
@@
We show that we can steer the output of Chameleon by intervening on the image patches using language tokens and analyzing how this affects the textual output. Focusing on the color setup, if the representation for a color is similar between visual and language inputs, we hypothesize that we can replace the image hidden states corresponding to one pure color image patch with the unembedding of the language token for another color, and mislead the model to “perceive” the new color when asked about the image color. Note that replacing the hidden state is a more invasive intervention than addition. However, there is a confounder: the intervened word may lexically bias the model to generate the same word, without performing reasoning that incorporates the new color, because the desired answer after intervening is equivalent to the intervention word itself (unlike in the previous cases). To control for this, we show two colors in one image and only intervene at the positions corresponding to one color: if the intervention unconditionally and lexically biases the generation to the new color, this effect would (incorrectly) affect both colors. [^8] We tested settings that require more sophisticated reasoning such as asking for a country flag with the two colors, or asking about spatial relationships of the colors. They seem to be beyond the capability of Chameleon-7B—even without interventions, the model cannot answer the questions correctly.
@@ 5 Discussion
Our experiments across diverse languages and modalities give empirical support to the semantic hub hypothesis; language models seem to make efficient use of model capacity and learn to represent semantically similar inputs from different modalities near one another. Such a semantic hub has been hypothesized to exist in the human brain (Patterson et al., 2007; Patterson & Ralph, 2016; Ralph et al., 2017). Inasmuch as both the human brain and language models are resource-constrained information processing systems, it is perhaps unsurprising that both systems make use of a semantic hub.
We note however that this strategy may not always be a desirable property of language models. Wendler et al. (2024) conjectured that language models might inherit biases present in the training data of the dominant language; if true, an extreme implementation of this strategy could force an alignment of ideology across languages and potentially harm inclusivity. Similarly, internalizing arithmetic expressions in natural language may not be ideal as a part of the “algorithm” with which a language model implements arithmetic, considering their structural differences and the vast linguistic variation in number expressions in their composition (c.f. Danish “tooghalvfems”; trans. ninety-two, but compositionally representing 2+(5-0.5)$\times$20) or even bases (c.f. the base-23 numeral system in the Kalam language (Laycock, 1975)). In general, an over-reliance on the dominant data type could be detrimental.
@@ 6 Related Work
A long line of work has investigated the representations of separately trained mono-data-type models, and showed that they can be aligned through a transformation. In the multilingual case, it has been found that separately trained word embeddings for different languages can be aligned (Mikolov et al., 2013; Smith et al., 2017; Cao et al., 2020; i.a.). Similarly, prior work has shown that visual representations and text representations from different models can be mapped together (Merullo et al., 2022; Koh et al., 2023; Maniparambil et al., 2024; i.a.). Lu et al. (2021) showed that language-only models can be minimally finetuned to achieve high performance in other modalities. Our work, in contrast, looks at a single static model that can process multiple types of input data, and finds that the resulting representations align, without needing a transformation. Concurrent work (Luo et al., 2024) found similar alignments of “task vectors” (Ilharco et al., 2022; Hendel et al., 2023) in vision-language models.
@@
Much past work has analyzed the representation evolution throughout transformer layers, inspecting how it affects reasoning (Yang et al., 2024), factuality (Chuang et al., 2024), knowledge (Jin et al., 2024), etc. From another angle, work on layer pruning and early exiting also speaks to the representation dynamics across layers (Gromov et al., 2024; Sanyal et al., 2024; i.a.). More mechanistically, Merullo et al. (2024), Todd et al. (2024), Hendel et al. (2023), i.a., try to more precisely characterize the representation changes algorithmically.

### v3

@@ 1 Introduction
To what extent is this idealized strategy adopted by actual models? Wendler et al. (2024) find that on simple synthetic tasks, Llama-2 (Touvron et al., 2023) maps various input languages into a shared “English space” before projecting back out into another language, hinting that it leverages this shared representation scheme to an extent, at least for different languages. We show that this is a much more general phenomenon: when a model processes inputs from multiple data types, there is a shared representation space, and this space is scaffolded by the LM’s inherently dominant language (usually English). By scaffolded, we mean that the shared space can be interpreted to an extent in the dominant data type via the logit lens (nostalgebraist, 2020). Following the cognitive science nomenclature, we call this shared representation space the LM’s “semantic hub.”
@@
Our work is complementary and distinct from prior work which finds structural similarities between the representation spaces of models trained (usually independently) on different data types, such as those showing that text representations from text-only LMs can be aligned, via a transformation, to vision/audio representations of modality-specific models (Ilharco et al., 2021; Merullo et al., 2022; Li et al., 2023; Ngo & Kim, 2024; Huh et al., 2024; i.a.), and the literature on cross-lingual word embedding alignment (Mikolov et al., 2013; Artetxe et al., 2017; Conneau et al., 2018; Schuster et al., 2019; i.a.). We instead show that an LM trained on multiple data types represents and processes them in a shared space without requiring explicit alignment transformation. We hope our findings shed light on ways to more easily interpret the mechanisms of current models and motivate future work aimed at better controlling models using these insights. [^2] We release our code at https://github.com/ZhaofengWu/semantic-hub.
@@ 2 The Semantic Hub Hypothesis
An LM parameterizes a similar process: it uses $M_{\text{LM}}$ to map various input data types into a representation space $\mathcal{S}_{\text{LM}}\subseteq\mathbb{R}^{d}$ (early layers), performs computations in the space (middle layers), and verbalizes the output via $V_{\text{LM}}$ (end layers and the LM head). However, it is unknown as to how different data types are structured in the representation space. For example, one possibility is that the LM partitions $\mathbb{R}^{d}$ into disjoint subspaces for each data type and processes them separately. We instead hypothesize that LMs, through training, learn to represent and process different data types in a shared representation space that functions as a modality-agnostic “semantic hub.” That is, semantically similar inputs $w_{1:t}^{z_{1}}$ and $w_{1:t^{\prime}}^{z_{2}}$ from distinct data types—for example texts in different languages that are mutual translations—are similarly mapped in $\mathcal{S}_{\text{LM}}$; informally, $M_{\text{LM}}(w_{1:t}^{z_{1}})\approx M_{\text{LM}}(w_{1:t^{\prime}}^{z_{2}})$. However, absolute similarity measures (i.e., $\operatorname{sim}(M_{\text{LM}}(w_{1:t}^{z_{1}}),M_{\text{LM}}(w_{1:t^{\prime}}^{z_{2}}))$) are generally difficult and unintuitive to interpret in high dimensional spaces. [^3] See for example Beyer et al. (1999). Most prior work on probing also implicitly uses relative similarity measures since the similarity scores are normalized over a finite label set. We thus focus on relative similarity measures, taking a semantically unrelated sequence $u_{1:t^{\prime}}^{z_{2}}$, and evaluating whether $w_{1:t}^{z_{1}}$ is closer to $w_{1:t^{\prime}}^{z_{2}}$ than $u_{1:t^{\prime\prime}}^{z_{2}}$, as illustrated on the left of Figure 2. Formally:
@@ 3.1 Multilingual
Translation datasets enable a direct test for Eq. 1, with semantically equivalent cross-lingual sentences as $w_{1:t}^{z_{1}}$ and $w_{1:t^{\prime}}^{z_{2}}$ and a randomly chosen non-matching sentence as $u_{1:t^{\prime\prime}}^{z_{2}}$. We use the professionally-translated English-Chinese parallel sentences from Chen et al. (2016) ($N=5260$). For each sentence pair, we use a template to transform each sentence and compute the representation cosine similarity for each layer, using the last token position as the sentence representation, which has been shown to preserve sentential information (Morris et al., 2023). We consider two English-dominant LMs, Llama-2 and Llama-3 (Llama-3-Team, 2024), one Chinese-dominant LM, Baichuan-2 (Yang et al., 2023), and one multilingual LM, BLOOM (BigScience, 2023), specifically the 7B/8B variants. See §A.1 for more details.
@@
Figure 4 shows the probability of English and Chinese (and of the top language when it is neither) for each layer on 10,000 English/Chinese Wikipedia sentences. When English-dominant models process Chinese text, Wendler et al.’s finding generalizes, where English dominates in the intermediate layers and Chinese only dominates in the final layers. On the Chinese-dominant LM, this trend flips: when processing English text, its intermediate layers are closer to the Chinese space and the final layers are closer to the English space. For BLOOM, a multilingual model with a relatively balanced training language mixture, we do not see a clear dominating language in the intermediate layers; when we manually inspect the closest token, in most cases we observe symbols with no clear semantics (though this does not mean it does not have a unified representation space: see Exp. 1). 70B model trends are also highly similar to the 7B/8B ones.
@@ 3.3 Code
Many recent LMs are trained on code corpora (Llama-3-Team, 2024; Gemini-Team, 2024; i.a.). We find that they similarly process code input by projecting it into a unified representation space shared with regular language tokens. Figure 7 shows some examples, where the LM in the intermediate layers tends to verbalize the future in free-form English, unconstrained by program syntax. E.g., in the first program, given the Python prefix “... for idx, e in enumerate(numbers): for idx2, e2 in enumerate(numbers”, instead of the ground-truth continuation in Python “): if idx != idx2:”, the most salient intermediate token is “except”, likely attempting to predict in English “(for each element in numbers) except if it is equal to idx”. Similarly, in a list literal expression “[1.0, 2.0,” instead of continuing in Python “ 3.0”, it predicts “and”, which is a natural way to continue in English. In these cases, it is difficult to obtain semantically equivalent English-Python prefix pairs like in §3.1 for testing Eq. 1555We may argue that functions and their specifications constitute such pairs, but we found that their correspondence is often too abstract and non-exact to manifest as similar representations., so we only test Eq. 3 across a few targeted cases in Python below.
@@
Function arguments have names in the definition, such as “range(start, end, step)”; but when invoked, they are filled with actual context-appropriate expressions. We call the argument names “semantic roles”, and the context-specific expressions the “surface forms”, inspired by thematic relations in linguistics (Fillmore, 1968; Jackendoff, 1974; i.a.). Like in the second example in Figure 1, we show that LMs predict the arguments by first “thinking” about their semantic role ($\tau^{z^{\star}}$) and then instantiating with surface-constrained expressions ($\tau^{z^{\circ}}$). We extract all function calls and arguments from MBPP with simple filtering, resulting in 540 arguments (see §A.2 for details). For each argument, we use the logit lens to inspect the hidden states $h_{t}^{\ell}$ at the preceding token (“(” or “,”). For each argument, Figure 8(b) visualizes if the semantic role or the surface form is closer to each layer’s hidden state. The semantic role ($\tau^{z^{\star}}$) dominates for the early to middle layers, and only in the final layers do the representations converge towards the surface form argument ($\tau^{z^{\circ}}$).
@@ 3.5 Visual Input
Though not constituting exact semantic equivalence, an image paired with its caption provides one possible test for Eq. 1. We take 1000 images and corresponding captions in the MSCOCO dataset (Lin et al., 2014) and measure their hidden states cosine similarity in LLaVA-7B (Liu et al., 2023) and Chameleon-7B (Chameleon-Team, 2024). Again, we subtract the average similarity between non-matching image-caption pairs as a baseline, separately for each layer. Figure 11(a) shows that semantically matching inputs across modalities are closer to one another than would be expected from chance, as in the translation experiments. Importantly, these models do not explicitly optimize cross-modal representation similarity (unlike in CLIP (Radford et al., 2021)), and such similarity only emerges through autoregressive training.
@@ 4 Intervening in the Semantic Hub
Past work has shown that (monolingual) interventions in the middle layers can steer the output of LMs in predictable ways (Subramani et al., 2022; Turner et al., 2024; Rimsky et al., 2024; i.a.). If the English-dominant LMs have a shared representation space, we should be able to intervene on this space in English even when processing other languages (Dumas et al., 2024). We use a popular hidden space intervention technique, Activation Addition (ActAdd; Turner et al., 2024), which works by: (1) taking a pair of contrasting steering words that semantically represent the steering effect (e.g., “Good” and “Bad” for sentiment steering), (2) taking their hidden state difference at an intermediate layer, and (3) scaling and adding the steering vector to the hidden states for the original forward pass of the regular generation process, at the same layer, at the beginning of the sequence. We generalize their sentiment-steering experiment cross-lingually. See Turner et al. (2024) for details.
@@
Based on our “semantic role” observation in §3.3, we intervene using the “range” function. We focus on two overloaded versions of “range”: “range(start, end)” and “range(end)”. If the semantic hub causally affects model output, then we can intervene in it to select the function version to use after “range(” in some context, using the English words “start” and “end”. Specifically, we take all single-argument “range(end)” calls in the MBPP dataset ($N=159$) and attempt to expand it into “range(0, end)”. We use a similar intervention method, except simply using the unembeddings of trigger tokens instead of intermediate LM hidden states. We use as trigger tokens (“start” – “end”) and add the intervention vector to the hidden states corresponding to the open parenthesis “(” at an intermediate layer (layer 17). For all these “range” calls in the dataset, we let Llama-2888We do not consider Llama-3 in this case because its default behavior usually generates “range(0, end)” in the first place, and it is unclear how to steer from “range(0, end)” to “range(end)”. generate without and with intervention. Figure 12(b) shows that, with increasing intervention strength, more instances are successfully steered to “range(0, end)”, up to 67%.
@@
We show that we can steer the output of Chameleon by intervening on the image patches using language tokens and analyze how this affects the textual output. Focusing on the color setup, if the representation for a color is similar between visual and language inputs, we hypothesize that we can replace the image hidden states corresponding to one pure color image patch with the unembedding of the language token for another color, and mislead the model to “perceive” the new color when asked about the image color. Note that replacing the hidden state is a more invasive intervention than addition. However, there is a confounder: the intervened word may lexically bias the model to generate the same word, without performing reasoning that incorporates the new color, because the desired answer after intervening is equivalent to the intervention word itself (unlike in the previous cases). To control for this, we show two colors in one image and only intervene at the positions corresponding to one color: if the intervention unconditionally and lexically biases the generation to the new color, this effect would (incorrectly) affect both colors. [^9] We tested settings that require more sophisticated reasoning such as asking for a country flag with the two colors, or asking about spatial relationships of the colors. They seem to be beyond the capability of Chameleon-7B—even without interventions, the model cannot answer the questions correctly.
@@ 5 Discussion
Our experiments across diverse languages and modalities give empirical support to the semantic hub hypothesis; language models seem to make efficient use of model capacity and learn to represent semantically similar inputs from different modalities near one another. Such a semantic hub has been hypothesized to exist in the human brain (Patterson et al., 2007; Patterson & Ralph, 2016; Ralph et al., 2017). Inasmuch as both the human brain and language models are resource-constrained information processing systems, it is perhaps unsurprising that both systems make use of a semantic hub. This also supports prior findings that multilingual models usually reason best in English (Shi et al., 2023).
We note however that this dominant-language-reliance may not always be a desirable property of language models. Wendler et al. (2024) conjectured that language models might inherit biases present in the training data of the dominant language; if true, an extreme implementation of this strategy could force an alignment of ideology across languages and potentially harm inclusivity. Similarly, internalizing arithmetic expressions in natural language may not be ideal as a part of the “algorithm” with which a language model implements arithmetic, considering their structural differences and the vast linguistic variation in number expressions in their composition (c.f. Danish “tooghalvfems”; trans. ninety-two, but compositionally representing 2+(5-0.5)$\times$20) or even bases (c.f. the base-23 numeral system in the Kalam language (Laycock, 1975)). In general, an over-reliance on the dominant data type could be detrimental.
@@ 6 Related Work
A long line of work has investigated the representations of separately trained mono-data-type models, and showed that they can be aligned through a transformation. In the multilingual case, it has been found that separately trained word embeddings for different languages can be aligned (Mikolov et al., 2013; Smith et al., 2017; Cao et al., 2020; i.a.). Similarly, prior work has shown that visual representations and text representations from different models can be mapped together (Merullo et al., 2022; Koh et al., 2023; Maniparambil et al., 2024; i.a.). Huh et al. (2024) argued that these are possible because the different data modalities are projections of the same underlying reality. Lu et al. (2021) showed that language-only models can be minimally finetuned to achieve high performance in other modalities. Our work, in contrast, looks at a single static model that processes multiple input data types, and finds that the resulting representations align, without needing a transformation. Concurrent work (Luo et al., 2024) found similar alignments of “task vectors” (Ilharco et al., 2022; Hendel et al., 2023) in vision-language models.
@@
Much past work has analyzed the representation evolution throughout transformer layers, inspecting how it affects reasoning (Yang et al., 2024), factuality (Chuang et al., 2024), knowledge (Jin et al., 2024), etc. From another angle, work on layer pruning and early exiting also speaks to the representation dynamics across layers (Gromov et al., 2024; Sanyal et al., 2024; i.a.). More mechanistically, Merullo et al. (2024), Todd et al. (2024), Hendel et al. (2023), i.a., more precisely characterized the representation changes algorithmically.